OpenStudy (anonymous):

If \(\frac{ x+3y }{ 2x-y }=\frac{ 2 }{ 3 }\) , then\(\frac{ x-y }{ x+y}\) =?

hero (hero):

\[\frac{x + 3y}{2x - y} = \frac{2}{3}\] 3(x + 3y) = 2(2x - y) 3x + 9y = 4x - 2y 2y + 9y = 4x - 3x (2 + 9)y = (4 - 3)x 11y = x \[\frac{11}{1} = \frac{x}{y}\] \[\frac{x - y}{x + y} = \frac{11 - 1}{11 + 1} = \frac{10}{12} = \frac{5}{6}\]

OpenStudy (anonymous):

ohh, thanks, i know how to do now:)

hero (hero):

Do you have another one?

hero (hero):

That one was convenient because we were able to isolate x relatively easy.

hero (hero):

These are not usually that straight forward.